Need molecular orbitals of 1,3-butadiene
The molecular orbitals of 1, 3-butadiene are given
Answer #1

1,3-butadiene has two pi bonds which corresponds to four pi electrons.

HOMO is Highest Occupied Molecular Orbital

LUMO is Lowest Unoccupied Molecular Orbital

Sapling Learning Март The molecular orbitals of 1,3-butadiene are given below. 1) Fill in the pi electrons expected in the gr
Answer #2
https://youtu.be/k3DyuXyeaFU
source: YouTube
answered by: Vinod Kumari
Answer #3

1,3-butadiene has two pi bonds which corresponds to four pi electrons.

HOMO is Highest Occupied Molecular Orbital

LUMO is Lowest Unoccupied Molecular Orbital

1)
An electron pair is present in \pi _1 orbital
An electron pair is present in \pi _2 orbital
\pi _3 and \pi _4 orbitals are empty
2) HOMO is \pi _2
LUMO is \pi _3


answered by: ANURANJAN SARSAM